OpenStudy (anonymous):

Evaluate the function for x = –2a, if a = 3. f(x) = –4x + 7

OpenStudy (anonymous):

someone please helllpppppp!

OpenStudy (helder_edwin):

if a=3 then \[\large x=-2a=(-2)(3)=-6 \] so since x=-6 \[\large f(-6)=-4(-6)+7=24+7=31 \]

OpenStudy (anonymous):

thank you helps soooo much :D I have another question... Which of the following does not describe the independent variable? A. It is the domain. B. It is the y-values. C. It is the input. D. It is the x-values.

OpenStudy (anonymous):

i believe it is D but im unsure

OpenStudy (helder_edwin):

the domain is the set of values the independent variable takes on. the x-values are the input of a fuction which is the independent variable.

OpenStudy (anonymous):

so the answer will be a A?

OpenStudy (helder_edwin):

no. A, C and D, describe the same thing

OpenStudy (anonymous):

Thank you :)

OpenStudy (helder_edwin):

u r welcome